OpenStudy (anonymous):

f(x) = 5x+3, then what is f^–1(x) ?

OpenStudy (anonymous):

your function says 1) multiply by 5 2) add 3 inverse will say 1) subtract 3 2) divide by 5

OpenStudy (anonymous):

you can write it as \[f^{-1}(x)=\frac{x-3}{5}\]

OpenStudy (anonymous):

fist replace f(x) with y. y=5x+3 now replace each y with x and each x with y x=5y+3 now solve for y

OpenStudy (anonymous):

x-3/5=y?

OpenStudy (anonymous):

yes now replace y with f^-1(x)

OpenStudy (anonymous):

does doing that change my answer or its just to write it right?

OpenStudy (anonymous):

its just to write it like that your answer remains same f^-(x)=(x-3)/5